Circular No. 5943

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

NOVA SAGITTARII 1994 A. C. Gilmore, provides the following photometry, obtained with the

0.6-m f/16 Cassegrain telescope at Mt. John University Observatory on Feb. 27.68 UT under marginal conditions, with the region low in a moonlit sky (uncertainty ±0.1 mag): V = 8.53, U − B = +0.97, B − V = +1.44, V − R = +0.84, V − I = +1.55. P. M. Kilmartin has measured the following precise position from two exposures by Gilmore with the 0.15-m f/15 astrograph at Mt. John on Mar. 2.62 UT: α = 18^{h}47^{m}37^{s}.29, δ = −21^{o}27′01′′.4 (equinox 1950.0).

Magnitude estimates: Feb. 13.85 UT, 8.9 (Y. Sakurai, Mito, Ibaragi, Japan, Fuji Super G400 Color film; prediscovery observation); 17.84, 8.7 (Sakurai); 22.83, 8.3 (Sakurai); 27.68, 8.7 (A. F. A. Jones, Nelson, New Zealand, visual); 27.81. 8.1 (Sakurai).

GRB 940301 On behalf of the BATSE/COMPTEL/NMSU Rapid Burst Response

Campaign (cf. IAUC 5937, but including now also C. Kouveliotou), R. M. Kippen reports: “COMPTEL has imaged a strong cosmic γ-ray burst that was first detected by the BATSE experiment on Mar. 1.84071 UT. Significant emission > 0.72 MeV, measured for > 40 s, consists of an intense single peak with substructure. The preliminary COMPTEL imaging analysis based on ∼ 170 telescope events yields a roughly circular region centered at α = 6^{h}49^{m}.7, δ = +63^{o}55′ (equinox 2000.0) with uncertainty radii of 1^{o}.0 (2σ-confidence) and 1^{o}.5 (3σ-confidence).”

COMET McNAUGHT-RUSSELL (1993v) S. Nakano, Sumoto, Japan, reports that he and I. Hasegawa have ex-

amined the possibility that this comet may be identical with the comet of 574, for which Hasegawa (1979, Publ. Astron. Soc. Japan 31, 257) has given a rather similar orbit. The orbit of 1993v on IAUC 5926 indicates a revolution period of ∼ 1440 years, and a recent refinement by Nakano gives 1430±30 years. He also finds that the two-month span of 1993–1994 observations can be satisfactorily linked with the positions deduced by Hasegawa for 574 Apr. 4, May 8 and 23, provided that one accepts residuals in 574 of up to 1^{o}.6.

1994 March 2 Brian G. Marsden